# Moving horizon estimation for nonlinear systems with time-varying parameters

### v1

@@ 1 Introduction
Notation   We denote the set of integers by $\mathbb{I}$, the set of all integers greater than or equal to $a\in\mathbb{I}$ by $\mathbb{I}_{\geq a}$, and the set of integers in the interval $[a,b]$ for any $a,b\in\mathbb{I}$ by $\mathbb{I}_{[a,b]}$. The $n\times n$ identity matrix is denoted by $I_{n}$ and the $n\times m$ zero matrix by $0_{n\times m}$. The weighted Euclidean norm of a vector $x\in\mathbb{R}^{n}$ with respect to a positive definite matrix $Q=Q^{\top}$ is defined as $\|x\|_{Q}=\sqrt{x^{\top}Qx}$ with $\|x\|=\|x\|_{I_{n}}$. The minimal and maximal eigenvalues of $Q$ are denoted by $\underline{\lambda}(Q)$ and $\overline{\lambda}(Q)$, respectively. For two matrices $A=A^{\top}$ and $B=B^{\top}$, we write $A\succeq B$ ($A\succ B$) if $A-B$ is positive semi-definite (positive definite). For $A,B\succ 0$, the maximum generalized eigenvalue (i.e., the largest scalar $\lambda$ satisfying $\det(A-\lambda B)=0$) is denoted by $\overline{\lambda}(A,B)$.
@@ 3 Detectability and observability concepts
$X{=}\left(\left\{(x_{t},z_{t},u_{t},w_{t})\right\}_{t=0}^{T-1},\left\{(\tilde{x}_{t},\tilde{z}_{t},u_{t},\tilde{w}_{t})\right\}_{t=0}^{T-1}\right)\in\mathbb{D}^{T}\times\mathbb{D}^{T}$
@@
[…]
@@ 4.2 Stability analysis
The following lemma establishes boundedness of the estimation error on $[t_{1},t]$ and a decrease in Lyapunov coordinates on each interval $[t_{m+1},t_{m}]$ for all $m\in\mathbb{I}_{[1,k]}$.
@@
[…]
@@
$\leq\mu^{k}\Gamma(c,\hat{x}_{l},x_{l},\bar{z}_{l},z_{l})+\sum_{m=1}^{k}\mu^{m-1}\left(c_{Q,2}\sum_{r={t_{m+1}}}^{t_{m}-1}\|w_{r}\|_{Q}\right)^{2},$ (36)

### v2

@@ Abstract
{textblock*}
(1.5cm,29.3cm) © 2024 The Authors. This work has been accepted to IFAC for publication under a Creative Commons Licence CC BY-NC-ND.
10.1016/j.ifacol.2024.09.053
@@ 1 Introduction
Notation   We denote the set of integers by $\mathbb{I}$, the set of all integers greater than or equal to $a\in\mathbb{I}$ by $\mathbb{I}_{\geq a}$, and the set of integers in the interval $[a,b]$ for any $a,b\in\mathbb{I}$ by $\mathbb{I}_{[a,b]}$. The $n\times n$ identity matrix is denoted by $I_{n}$ and the $n\times m$ zero matrix by $0_{n\times m}$. The weighted Euclidean norm of a vector $x\in\mathbb{R}^{n}$ with respect to a positive definite matrix $Q=Q^{\top}$ is defined as $\|x\|_{Q}=\sqrt{x^{\top}Qx}$ with $\|x\|=\|x\|_{I_{n}}$. The minimal and maximal eigenvalues of $Q$ are denoted by $\underline{\lambda}(Q)$ and $\overline{\lambda}(Q)$, respectively. For two matrices $A=A^{\top}$ and $B=B^{\top}$, we write $A\succeq B$ ($A\succ B$) if ABA-B is positive semi-definite (positive definite). For $A,B\succ 0$, the maximum generalized eigenvalue (i.e., the largest scalar $\lambda$ satisfying $\det(A-\lambda B)=0$) is denoted by $\overline{\lambda}(A,B)$.
@@ 3 Detectability and observability concepts
$X{=}\left(\hskip-2.15277pt\left\{(x_{t},z_{t},u_{t},w_{t})\right\}_{t=0}^{T-1},\left\{(\tilde{x}_{t},\tilde{z}_{t},u_{t},\tilde{w}_{t})\right\}_{t=0}^{T-1}\right)\in\mathbb{D}^{T}\times\mathbb{D}^{T}$
@@
$\mathbb{E}_{T}:=\Big{\{}$ $\left(\left\{(x_{t},z_{t},u_{t},w_{t})\right\}_{t=0}^{T-1},\left\{(\tilde{x}_{t},\tilde{z}_{t},u_{t},\tilde{w}_{t})\right\}_{t=0}^{T-1}\right)$
@@ 4.2 Stability analysis
The following lemma establishes boundedness of the estimation error on $t_{1},t]$ and a decrease in Lyapunov coordinates on each interval $t_{m+1},t_{m}]$ for all $m\in\mathbb{I}_{[1,k]}$.
@@
$c_{\mathrm{Q},2}:=\max\Bigg{\{}$ $\mu\left(\frac{c_{1}^{\prime}\max\{2\overline{\lambda}(\overline{V},\underline{V}),\eta^{-N}\}}{1-\rho}+2\right),$
@@
$\leq\mu^{k}\Gamma(c,\hat{x}_{l},x_{l},\bar{z}_{l},z_{l})+\sum_{m=1}^{k}\mu^{m-1}\left(c_{Q,2}\sum_{r={t_{m+1}}}^{t_{m}-1}\hskip-4.30554pt\|w_{r}\|_{Q}\right)^{2},$ (36)